# GestureHYDRA: Semantic Co-speech Gesture Synthesis via Hybrid Modality Diffusion Transformer and Cascaded-Synchronized Retrieval-Augmented Generation

### v1

@@ 4.3 Evaluation Results
On the unseen test set, our method also outperforms existing SOTA methods on all metrics, demonstrating its superior generalization capability. In contrast, the VQ-VAE-based counterparts, suffer from limited generalization due to their limited codebooks. In summary, our system can generate motions that reach a realistic level, with robust performance in rhythm alignment and semantic gesture generation quality on both seen and unseen data.
@@ 4.4 Ablation study
To clearly demonstrate the effectiveness of “3D kp loss”, we additionally perform a study on a downstream task (i.e., human video generation), which produces corresponding human videos based on our generated co-speech gestures by using a human video generation model [16]. As shown in Fig. 5, our full model is able to synthesize a stable human body with satisfying hand-to-table interactions instead of floating elbows as in the results of “w/o 3D kp loss”, which further facilitates the effectiveness of our $L_{kp}$.
@@ A.1 Dataset Details
After obtaining the clips, we follow SHOW [51] to reconstruct the SMPL-X [32] parameters from monocular videos. Our work focuses on body and gesture generation, so we do not reconstruct facial expression coefficients. Additionally, while SHOW uses PyMAF-X [54] to initialize hand poses, we replace PyMAF-X with the more powerful hamer [33] to improve the quality of initial hand reconstruction.
@@ A.2 Automatic Annotation Approach
Our framework is also equipped with an automatic semantic gesture annotation procedure (shown in Fig. 6) that can extract semantic gestures from a video of each target person. Specifically, we first prepare 18 pre-defined standard semantic gestures in the form of 2D skeletons by using DWPose [50] detection, which can be regarded as semantic gesture templates. Given a target video, we first use automatic-speech-recognition (ASR) technique to convert its audio into text, and then locate the segments aligned with special trigger words (e.g., ”first”, ”left” , and ”hello”). By performing DWPose detection on these segments, we are able to calculate the similarity between the detected gestures and the corresponding pre-defined gesture template. It’s worth noting that semantic gestures may not be triggered every time when a person speaks the same word, thus such a process is necessary to filter relevant segments for personalized gesture collection. Once a segment exhibits high consistency with the gesture template, we continue to perform the 3D human body reconstruction process, and the reconstructed result is accepted as a standard example of the corresponding gesture. Normally, we collect at least two examples for each gesture unless a specific gesture is quite rare or missing in the captured data.

### v2

@@ 4.3 Evaluation Results
On the unseen test set, our method also outperforms existing SOTA methods on all metrics, demonstrating its superior generalization capability. In contrast, the VQ-VAE-based counterparts suffer from limited generalization due to their limited codebooks. In summary, our system can generate motions that reach a realistic level, with robust performance in rhythm alignment and semantic gesture generation quality on both seen and unseen data.
@@ 4.4 Ablation study
To clearly demonstrate the effectiveness of “3D kp loss”, we additionally perform a study on a downstream task (i.e., human video generation), which produces corresponding human videos based on our generated co-speech gestures by using a human video generation model [animateanyone]. As shown in Fig. 5, our full model is able to synthesize a stable human body with satisfying hand-to-table interactions instead of floating elbows as in the results of “w/o 3D kp loss”, which further facilitates the effectiveness of our $L_{kp}$.
@@ A.1 Dataset Details
After obtaining the clips, we follow SHOW [talkshow] to reconstruct the SMPL-X [smplx] parameters from monocular videos. Our work focuses on body and gesture generation, so we do not reconstruct facial expression coefficients. Additionally, while SHOW uses PyMAF-X [PyMAF] to initialize hand poses, we replace PyMAF-X with the more powerful hamer [hamer] to improve the quality of initial hand reconstruction.
@@ A.2 Automatic Annotation Approach
Our framework is also equipped with an automatic semantic gesture annotation procedure (shown in Fig. 6) that can extract semantic gestures from a video of each target person. Specifically, we first prepare 18 pre-defined standard semantic gestures in the form of 2D skeletons by using DWPose [dwpose] detection, which can be regarded as semantic gesture templates. Given a target video, we first use automatic-speech-recognition (ASR) technique to convert its audio into text, and then locate the segments aligned with special trigger words (e.g., ”first”, ”left” , and ”hello”). By performing DWPose detection on these segments, we are able to calculate the similarity between the detected gestures and the corresponding pre-defined gesture template. It’s worth noting that semantic gestures may not be triggered every time when a person speaks the same word, thus such a process is necessary to filter relevant segments for personalized gesture collection. Once a segment exhibits high consistency with the gesture template, we continue to perform the 3D human body reconstruction process, and the reconstructed result is accepted as a standard example of the corresponding gesture. Normally, we collect at least two examples for each gesture unless a specific gesture is quite rare or missing in the captured data.